# Simulation of prosthetic vision with PRIMA system and enhancement of face representation

### v2

@@ Abstract
Objective Patients implanted with the PRIMA photovoltaic subretinal prosthesis in geographic atrophy report form vision with the average acuity matching the 100m pixel size. Although this remarkable outcome enables them to read and write, they report difficulty with perceiving faces. Despite the pixelated stimulation, patients report seeing smooth patterns rather than dots. This paper provides a novel, non-pixelated algorithm for simulating prosthetic vision the way it is experienced by PRIMA patients, compares the algorithm’s predictions to clinical perceptual outcomes, and offers computer vision and machine learning (ML) methods to improve face representation.
Approach Our simulation algorithm integrates a grayscale filter, spatial resolution filter, and contrast filter. This accounts for the limited sampling density of the retinal implant (pixel pitch), as well as the reduced contrast sensitivity of prosthetic vision. Patterns of Landolt C and faces created using this simulation algorithm are compared to reports from actual PRIMA users. To recover the facial features lost in prosthetic vision due to limited resolution or contrast, we apply an ML facial landmarking model as well as contrast adjusting tone curves to the face image prior to its projection onto the photovoltaic retinal implant.
Main results Prosthetic vision simulated using the above algorithm matches the maximum letter acuity observed in clinical studies as well as patients’ subjective descriptions of perceived facial features. Application of the inversed contrast filter to the image prior to its projection onto the implant helps preserve the contrast in prosthetic vision. Identification of the facial features using an ML facial landmarking model and accentuating them using line widths larger than the resolution limit of the implant further improve the face representation in prosthetic vision.
Significance Spatial and contrast constraints of prosthetic vision limit resolvable features and degrade natural images. ML based methods and contrast adjustments prior to image projection onto the implant mitigate some limitations and improve face representation. We also show that even though higher spatial resolution can be expected with implants having smaller pixels, contrast enhancement still remains essential for face recognition.
Jungyeon Park1†∗, Anna Kochnev Goldstein2∗, Yueming Zhuo2,
Nathan Jensen2, and Daniel Palanker1,3
1Hansen Experimental Physics Laboratory, Stanford University, Stanford, CA, USA
2Department of Electrical Engineering, Stanford University, Stanford, CA, USA
3Department of Ophthalmology, Stanford University, Stanford, CA, USA
†Corresponding author: jungyeon@stanford.edu
*These authors contributed equally.
Keywords: Prosthetic Vision, Retinal Implant, PRIMA, AMD, Face Perception, Face Recognition, Image Processing, Machine Learning
@@ 1 Introduction
Retinal degenerative diseases resulting in the progressive loss of photoreceptors, such as Retinitis Pigmentosa, Stargardt Disease and Age-related Macular Degeneration (AMD), are among the leading causes of incurable blindness today [1]. While photoreceptors are lost in retinal degeneration, the inner retinal neurons survive to a large extent [2-4]. Retinal prostheses are designed to reintroduce information into the visual system by electrically stimulating these remaining retinal neurons.
@@
PRIMA (Science Corp., Alameda, CA) is a photovoltaic subretinal prosthesis developed to restore vision in the geographic atrophy (GA) region of AMD patients. The 2×2mm PRIMA implant is 30m thick and is composed of 378 tiled pixels, each 100m wide, containing active and return electrodes. Each pixel in this implant converts light into electric current to stimulate the second-order retinal neurons, primarily the bipolar cells [5]. Because photovoltaic pixels require light much brighter than ambient for stimulation and electric current should be pulsed to preserve the charge balance, images captured by a camera are projected from augmented-reality (AR) glasses onto the implant using pulsed intensified light (Figure 1A). To avoid any confounding effects of this intense light on the remaining photoreceptors outside the GA, its wavelength is invisible—in the near-infrared part of the spectrum (880nm). A controller with a pocket computer provides image processing for enhancement of prosthetic vision, while the transparent glasses leave the peripheral natural vision unaffected. Since the photovoltaic implant is powered by light, it has no wires that could affect surrounding residual photoreceptors.
@@
In a series of two clinical trials, PRIMA was implanted in 43 AMD patients suffering from central vision loss due to GA. The implant enabled patients to reliably recognize letters and words, with acuity closely matching the maximum spatial resolution allowed by the 100m pixels [6, 7], which is 20/417. Normal vision of 20/20 corresponds to an angular resolution of 30 cycles per degree, and one degree in a human eye corresponds to 288m on the retina [8]. Since at least two pixels are required to resolve a cycle, the corresponding equivalent pixel size for normal vision is 288/30/2=4.8m. Therefore, 100m pixels correspond to acuity about 21 times (100/4.8) worse than normal. Using electronic zoom on the controller, patients improved the resolvable letter size, on average, by 23 ETDRS letters (4.6 lines) compared to the baseline, demonstrating clinically meaningful benefits of prosthetic central vision [9-11]. At the same time, patients reported a subjective experience of being able to identify big facial features, such as hairline, but unable to perceive finer facial details or identify faces. As such, facial perception remains one of their most desired features of prosthetic vision.
@@
To understand whether the facial perception obtained with prosthetic vision can be improved, one needs to envision what patients see when they look at a face using prosthetic vision. Several previous studies have offered illustrations of prosthetic vision, characterizing it as patterns of bright spots in a pixelated image [12-15]. However, PRIMA patients describe their precepts as smooth continuous white lines rather than a collection of round spots, even though the electric potential generated by the implant is indeed pixelated (Fig. 1B). It is unclear what part of the visual pathway converts the pixelated stimulation pattern applied to the bipolar cells into perception of smooth lines. It may be related to the perceptual completion phenomenon, where we tend to perceive visual features that aren’t present in the scene to reconcile it with familiar or expected patterns and objects [16].
@@
This paper offers a novel approach to simulation of prosthetic vision based on both objective clinical measures and subjective accounts of the PRIMA patients’ experience that aligns with the smooth pattern perception. In addition, based on this approach, it describes a pathway to improve facial perception using computer vision. The proposed methods hold the potential to provide enhanced face perception to existing PRIMA patients via a software update without any changes to the implanted hardware. In addition, we show the expected extent of improvements with the next-generation implants containing smaller and denser pixels.
@@ 2 Simulating the prosthetic vision of PRIMA patients
Retinal degeneration leads to progressive loss of contrast sensitivity and spatial resolution [17]. In prosthetic vision, the density (pitch) of the implant pixels imposes an upper limit on the spatial frequencies that can be sampled and transferred by the implant. Patients with subretinal implants describe their prosthetic vision as monochromatic (variations of white) [7, 18], with up to seven distinguishable contrast levels [19]. Prosthetic contrast sensitivity in rodents was also found to be about 12%, which corresponds to 8 distinguishable grey levels [20].
@@ 2.1 Resolution Reduction
The PRIMA implant is a 2x2mm photovoltaic array with 100m pixels tiled in an approximate 20x20 grid. Since at least two pixels are required to sample a cycle, the current device imposes a high frequency cutoff of 10 cycles per image. Therefore, we implemented a radial low-pass filter of the Fourier transformed image using a Tukey window with a radius of 10 cycles (Figure 2). Repeated sampling of the image with saccadic eye movements helps improve resolution beyond the sampling limit by as much as 30% [21-23], presumably based on similar principles to the super-resolution algorithms implemented with conventional cameras [24]. Therefore, to account for the improved acuity associated with microsaccades, we added a 30% apodization to the low-pass filter, ensuring that frequencies are gradually attenuated beyond the cutoff range with a tapering factor of 0.3.
@@ 2.2 Contrast Reduction
Clinical observations demonstrated the reduced contrast sensitivity in retinal degeneration and in prosthetic vision, but it is unclear whether that loss results in a more uniformly grey or a more distinctly black-and-white perception. Therefore, we simulated contrast sensitivity distortion by two transformations: (a) Gamma tone curve that converts the original image to a more uniformly grey, and (b) Sigmoid tone curve that results in a more distinctly black-and-white perception (Figure 3). Both curves retain the black (zero brightness) and white (maximum brightness) levels.
@@
2.3 Simulated Letters
To ensure that results of our simulation align with the acuity results reported in the clinical trials, we applied the algorithm to a grid of Landolt C optotypes of several sizes (Figure 5A). In this task, visual acuity is determined by the patient’s ability to identify the orientation of the gap in the letter C. As can be seen Figure 5 C, D, a gap of 0.8 pixels and larger can be resolved with both types of contrast transformations, matching the clinical results in the best PRIMA patients.
@@ 2.4 Simulated Faces
To simulate the perception of faces in prosthetic vision, the same algorithm of converting the original color image to grayscale, reducing the resolution to the PRIMA sampling limit, and compressing the contrast curve is applied. The simulated outputs are shown in Figure 6. The ringing effect in the output images, particularly visible at the top row, is a computational artifact of the Fourier filtering procedure and is not what patients experience. In principle, this ringing can be suppressed using a more complex filter design. However, since a long-tailed filter that includes higher frequencies cannot be easily justified as accurately representing the physical sampling of the device, we decided to keep this artifact in favor of a more straightforward filtering procedure. As can be seen in Figure 6, using prosthetic vision, one may be able to tell if they are looking at a face and identify the person’s hairline, but they don’t recognize the individual or identify facial expressions or emotions.
When several PRIMA patients were asked about their face perception using the device, they described the faces as a “cloud white” or “completely blurred.” Some reported that the hairline and the hair length were visible, and that they can partially see where the face ends and the background begins. One patient reported being able to partially see the nose and the lips. Although it remains unclear which one of the contrast reduction transformations is the most accurate for prosthetic vision with PRIMA, the illustrations in the two right columns (created using the sigmoid transform) stand in good agreement with these subjective descriptions of face perception obtained from the patients.
@@ 3.1 Improving Contrast Sensitivity
If the contrast reduction function of prosthetic vision is known, we could minimize its effect by applying the inverse function before projecting the image into the eye. Figure 7 shows the results of applying the inverse of the Gamma and the Sigmoid curves, respectively, to the original image before passing the resulting image through our prosthetic vision simulator. As one can see in comparison of the “original” columns to “inverse” ones, transforming the contrast of the original image in a way that counters the loss caused by prosthetic vision restores additional levels of gray. This allows us to better represent some important facial landmarks, such as eye sockets, mouth and cheek bones.
One limitation in such inversion is the fact that DMD projector in PRIMA glasses provides only 14 discrete levels of stimulation (adjusted by micromirrors ON time from 0.7 to 9.8ms in 0.7ms increments), which limits the fidelity of any "inversed" transformation. However, this hardware limitation might be alleviated in the future. More importantly, in our simulations, we know the exact function used to simulate the loss of contrast sensitivity. Hence, it is straightforward to apply the exact inverse of that function before the simulation algorithm and get the result shown in Figure 7. However, we don’t know the exact function governing the contrast loss in PRIMA patients. Therefore, clinical testing with individual patients is required to better understand the contrast loss function in order to design an inverse function that can improve contrast in prosthetic vision.
@@ 3.2 Reintroducing facial features
Spatial and contrast filters of simulated prosthetic vision with PRIMA remove most of the finer facial details, leaving just the face outline and hair. We can improve this outcome by making the fine facial features more pronounced in the original image before applying the simulation algorithm. To achieve this, we first identify the facial landmarks using the open-source, real-time “MediaPipe Face Landmarker” machine learning model from Google on the original color image. Then, we thicken the contours of the obtained key facial features—eyebrows, irises, and lips—to enable their passage through the spatial filters. The results of this procedure are shown in Figure 8.
To a normally sighted individual, the thicker features appear unnatural, but after the filters of prosthetic vision, the key facial elements reappear attenuated and closer to natural size. This should enable the patients to perceive a smile or establish eye contact—critical aspects of human communication and connection. This enhancement is purely computational and can run in real-time at video rate (30 fps), so it can be provided to existing patients as a software update without requiring any hardware modifications.
@@ 3.3 Face Perception with Future Implants
The 100m pixels in PRIMA impose a hard limit on the spatial frequencies that can be perceived by patients. As was shown with rats, implants with smaller pixels should provide higher resolution [25]. To illustrate the potential benefits of smaller pixels for face perception, we simulated prosthetic vision from 2x2mm implants composed of 50m and 20m pixels, with corresponding cutoff frequencies of 20 and 50 cycles per image. As shown in Figure 9, smaller pixels enable significantly higher spatial frequencies, allowing the finer details of the human face, such as the eyes, nose, eyebrows and lips, to be sampled by the device and visualized. However, despite the improvement in spatial resolution, the reduced contrast of prosthetic vision still makes the simulated image harder to perceive compared to original. Applying inversed contrast filters prior to the simulation algorithm optimizes the advantage of smaller implant pixels and should be sufficient for face recognition even without the ML-based enhancement of facial features.

### v3

# Simulation of prosthetic vision with the PRIMA system and enhancement of face representation

@@ Abstract
Objective. Patients implanted with the PRIMA photovoltaic subretinal prosthesis in geographic atrophy report form vision with the average acuity matching the 100m pixel size. Although this remarkable outcome enables them to read and write, they report difficulty with perceiving faces. Despite the pixelated stimulation, patients report seeing smooth patterns rather than dots. This paper provides a novel, non-pixelated algorithm for simulating prosthetic vision the way it is experienced by PRIMA patients, compares the algorithm’s predictions to clinical perceptual outcomes, and offers computer vision and machine learning (ML) methods to improve face representation.
Approach. Our simulation algorithm (ProViSim) integrates a grayscale filter, spatial resolution filter, and contrast filter. This accounts for the limited sampling density of the retinal implant (pixel pitch), as well as the reduced contrast sensitivity of prosthetic vision. Patterns of Landolt C and faces created using this simulator are compared to reports from actual PRIMA users. To recover the facial features lost in prosthetic vision due to limited resolution or contrast, we apply an ML facial landmarking model, as well as contrast-adjusting tone curves to the image prior to its projection onto the photovoltaic retinal implant.
Main results. Prosthetic vision simulated using the above algorithm matches the maximum letter acuity observed in clinical studies, as well as the patients’ subjective descriptions of perceived facial features. Applying the inversed contrast filter to the image prior to its projection onto the implant and accentuating the facial features using an ML facial landmarking model helps preserve the contrast in prosthetic vision, improves emotion recognition and reduces the response time.
Significance. Spatial and contrast constraints of prosthetic vision limit resolvable features and degrade natural images. ML based methods and contrast adjustments prior to image projection onto the implant mitigate some limitations and improve face representation. Even though higher spatial resolution can be expected with implants having smaller pixels, contrast enhancement still remains essential for face recognition.
Keywords: Prosthetic Vision, Retinal Implant, PRIMA, Age-related macular degeneration (AMD), Face Perception, Face Recognition, Image Processing, Machine Learning (ML), Facial Landmarking, Contrast Sensitivity
@@ 1 Introduction
Retinal degenerative diseases resulting in the progressive loss of photoreceptors, such as Retinitis Pigmentosa, Stargardt Disease and Age-related Macular Degeneration (AMD), are among the leading causes of incurable blindness today (smith_risk_2001). While photoreceptors are lost in retinal degeneration, the inner retinal neurons survive to a large extent (mazzoni_retinal_2008, humayun_morphometric_1999, kim_morphometric_2002). Retinal prostheses are designed to reintroduce information into the visual system by electrically stimulating these remaining retinal neurons.
@@
PRIMA (Science Corp., Alameda, CA) is a photovoltaic subretinal prosthesis developed to restore vision in the geographic atrophy (GA) region of AMD patients. The 2×2mm PRIMA implant is 30 thick and is composed of 378 hexagonally tiled pixels, each 100m wide, containing active and return electrodes. Each pixel in this implant converts light into electric current to stimulate the second-order retinal neurons, primarily the bipolar cells (lorach_photovoltaic_2015). Because photovoltaic pixels require light much brighter than ambient for stimulation and the electric current should be pulsed to preserve the charge balance, images captured by a camera are projected from augmented-reality glasses onto the implant using pulsed intensified light (Figure 1A). To avoid any confounding effects of this intense light on the remaining photoreceptors outside the GA, its wavelength is invisible - in the near-infrared part of the spectrum (880nm). A controller with a pocket computer provides image processing for enhancement of prosthetic vision, while the transparent glasses leave the peripheral natural vision unaffected. Since the photovoltaic implant is powered by light, it has no wires that could affect the surrounding residual photoreceptors.
@@
In a series of two clinical trials, PRIMA was implanted in 43 AMD patients suffering from central vision loss due to GA. The implant enabled patients to reliably recognize letters and words, with acuity closely matching the maximum spatial resolution allowed by the 100m pixels (palanker_photovoltaic_2020, palanker_simultaneous_2022), which is 20/417. Normal vision of 20/20 corresponds to an angular resolution of 30 cycles per degree, and one degree in a human eye corresponds to 288m on the retina (drasdo_non-linear_1974). Since at least two pixels are required to resolve a cycle, the corresponding equivalent pixel size for normal vision is 288/30/2=4.8m. Therefore, 100m pixels correspond to acuity about 21 times (100/4.8) worse than normal. Using electronic zoom on the controller, patients improved the resolvable letter size, on average, by 25 ETDRS letters (5 lines) compared to the baseline, demonstrating clinically meaningful benefits of prosthetic central vision (holz_subretinal_2025). At the same time, patients reported a subjective experience of being able to identify big facial features, such as the hairline, but unable to perceive finer facial details or identify faces. As such, facial perception remains one of their most desired features of prosthetic vision.
@@
To understand whether the facial perception obtained with prosthetic vision can be improved, one needs to envision what patients see when they look at a face using prosthetic vision. Several previous studies have offered illustrations of prosthetic vision, characterizing it as patterns of bright spots in a pixelated image (chen_simulating_2009, wang_sketch_2024, dagnelie_simulations_2011, moneer_enhancing_2023, chang_facial_2012). However, PRIMA patients describe their precepts as smooth continuous white lines rather than a collection of spots, even though the electric potential generated by the implant is indeed pixelated (Fig. 1B). It is unclear what part of the visual pathway converts the pixelated stimulation pattern applied to the bipolar cells into perception of smooth lines. It may be related to the combination of eye movements with the perceptual completion phenomenon, where we tend to perceive visual features that aren’t present in the scene to reconcile them with familiar or expected patterns and objects (komatsu_neural_2006). Regardless of the reason, existing simulators that produce pixelated images do not seem to be the right framework to understand prosthetic vision with PRIMA. Other studies suggest substitutes such as simplifying the original detailed visual scene based on the task at hand (beyeler_towards_2022) or modifications such as caricaturing (irons_face_2017).
@@
This paper offers a novel simulator of prosthetic vision (ProViSim) based on both objective clinical measures and subjective accounts of the PRIMA patients’ experience that aligns with their smooth pattern perception. In addition, based on the results produced by this simulator, it offers a pathway to improve facial perception using computer vision aimed to help patients recognize faces in the most natural way without relying on scene augmentation. The proposed enhancements, tested on 16 sighted individuals, hold the potential to provide enhanced face perception to existing PRIMA patients via a software update without any changes to the implanted hardware. In addition, we show the expected extent of improvements with the next-generation implants containing smaller pixels.
@@ 2 Simulating the Prosthetic Vision of PRIMA Patients
Retinal degeneration leads to progressive loss of contrast sensitivity and spatial resolution (vingopoulos_measuring_2021). In prosthetic vision, the density (pitch) of the implant pixels imposes an upper limit on the spatial frequencies that can be sampled and transferred by the implant. Patients with subretinal implants describe their prosthetic vision as monochromatic (variations of white) (palanker_photovoltaic_2020, palanker_simultaneous_2022), with up to seven distinguishable contrast levels (zrenner_subretinal_2011). Prosthetic contrast sensitivity in rodents was also found to be about 12%, corresponding to 8 distinguishable grey levels (ho_temporal_2018).
@@ 2.1 Resolution Reduction
The PRIMA implant is a 2x2mm photovoltaic array with 100m pixels tiled in an approximate 20x20 grid. Since at least two pixels are required to sample a cycle, the current device imposes a high frequency cutoff of 10 cycles per image. Therefore, we implemented a radial low-pass filter of the Fourier-transformed image using a Tukey window with a radius of 10 cycles (Figure 2). Repeated sampling of the image with saccadic eye movements helps improve resolution beyond the sampling limit by as much as 30% (abraham_active_2019, ratnam_fixational_2015, ratnam_benefits_2017), presumably based on similar principles to the super-resolution algorithms implemented with conventional cameras (ur_improved_1992). Therefore, to account for the improved acuity associated with microsaccades, we added a 30% apodization to the low-pass filter, ensuring that frequencies are gradually attenuated beyond the cutoff range with a tapering factor of 0.3.
@@ 2.2 Contrast Reduction
Clinical observations demonstrated the reduced contrast sensitivity in retinal degeneration and in prosthetic vision, but it is unclear whether that loss results in a more uniformly grey or a more distinctly black-and-white perception. Therefore, we simulated contrast sensitivity distortion by two transformations: (a) A Gamma tone curve that converts the original image to a more uniformly grey, and (b) a Sigmoid tone curve that results in a more distinctly black-and-white perception (Figure 3). While the Gamma transform retains normal black (zero brightness) and white (maximum brightness) levels, the Sigmoid transform also includes a horizontal offset parameter that may represent a stimulation threshold and also enables to adjust the black level such that the darkest level is not as dark as normal black.
@@
2.3 Testing on Letters
To ensure that the results of our simulation align with the acuity measurements in the clinical trials, we applied the algorithm to a grid of Landolt C optotypes of several sizes (Figure 5A). In this task, visual acuity is determined by the patient’s ability to identify the orientation of the gap in the letter C. As can be seen in Figure 5C, D, a gap of 0.8 pixels and larger can be resolved with both types of contrast transformations, matching the clinical results in the best PRIMA patients.
@@ 2.4 Simulated Faces
To simulate the perception of faces in prosthetic vision, the same algorithm of converting the original color image to grayscale, reducing the resolution to the PRIMA sampling limit, and compressing the contrast curve is applied. The simulated outputs are shown in Figure 6. The ringing effect in the output images, particularly visible at the top row, is a computational artifact of the Fourier filtering procedure and is not what patients experience. In principle, this ringing can be suppressed using a more complex filter design. However, since a long-tailed filter that includes higher frequencies cannot be easily justified as accurately representing the physical sampling of the device, we decided to keep this artifact in favor of a more straightforward filtering procedure. As can be seen in Figure 6, using prosthetic vision, one may be able to tell if they are looking at a face and identify the person’s hairline, but not recognize the individual or identify facial expressions or emotions.
Most PRIMA patients did not look at faces as this wasn’t the focus of the trial and they were not directly instructed to do so. However, three patients that did look at faces described them as “cloud white” or “completely blurred” and reported that they could partially see where the face ends and the background begin, the outline of the face, and the presence of the hair and nose. One patient could also recognize the hair outline, hair style, facing direction, and even partially see the eyes and lips. Although it remains unclear which one of the contrast reduction transformations is the most accurate for prosthetic vision with PRIMA, the illustrations in the two right columns of the sigmoid transform stand in good agreement with these subjective descriptions of face perception obtained from the patients.
@@ 3.1 Improving Contrast Sensitivity
If the contrast reduction function of prosthetic vision is known, we could minimize its effect by applying the inverse function before projecting the image into the eye. Figure 7 shows the results of applying the inverse of the Gamma and the Sigmoid curves, respectively, to the original image before passing the resulting image through our prosthetic vision simulator. Comparison of the “original” columns to the “inverse” ones demonstrates that transforming the contrast of the original image in a way that counters the loss caused by prosthetic vision restores some additional levels of gray. This allows better representing some important facial planes, such as the eye sockets, forehead, and cheek bones.
One limitation in such an inversion is the fact that the DMD projector in the PRIMA glasses provides only 14 discrete levels of stimulation (adjusted by the micromirrors’ ON time from 0.7 to 9.8ms in 0.7ms increments), which limits the fidelity of any "inversed" transformation. However, this hardware limitation might be alleviated in the future. More importantly, since in our simulations, we know the exact function used to simulate the loss of contrast sensitivity, it is straightforward to apply the exact inverse of that function prior to the simulation algorithm and get the result shown in Figure 7. However, since we don’t know yet the exact function governing the contrast loss in PRIMA patients, clinical testing is required to assess the contrast loss in individual patients before applying an inverse function.
@@ 3.2 Reintroducing Facial Features
Spatial and contrast filters of simulated prosthetic vision with PRIMA remove most of the finer facial details, leaving just the face outline and hair. We can improve this outcome by making the fine facial features more pronounced in color or thickness in the original image before applying the simulation algorithm. To achieve this, we first identify the facial landmarks using the open-source, real-time “MediaPipe Face Landmarker” machine learning model from Google on the original color image. Then, we can thicken and darken the contours of the obtained key facial features - eyebrows, irises, and lips - to enable their passage through the filters. To a normally sighted individual, the bolder features appear unnatural, but after the filters of prosthetic vision, the key facial elements reappear attenuated and closer to natural size. The results of this procedure are shown in Figure 8.
Figure 9 shows three other people expressing different emotions – disgust, surprise, and happiness, enhanced with landmarking, as well as landmarking and contrast combined. Without any enhancements, one can hardly identify any of the emotions. However, with preemptive facial landmarking, and even more so with both landmarking and contrast correction, the emotions become much easier to identify and the direction of gaze more pronounced. This should enable the patients to perceive a smile and establish eye contact - critical aspects of human communication and connection. As this enhancement is purely computational and can run in real-time at video rate (30 fps), it can be provided to existing patients as a software update without requiring any hardware modifications.
Figure 9 also illustrates the difference between light and dark skin tones. While lighter skin tones lose most inner facial features when processed by the simulator, darker skin tones maintain much more useful information. In fact, they can benefit from slightly accentuating the facial features with facial landmarking, but applying contrast correction makes them unresolvedly dark. Therefore, no contrast correction was applied on the bottom row. Darker hair colors also remain more visible and help obtain the outline of the face and separate it from the background. The relative contrast between the face and the hair is also significant.
As contrast correction already restores a lot of the information in faces with lighter skin tones, the landmarking applied on these images was much gentler in both thickness and color. In practice, the thickness of the drawn features, their color, and the application of contrast correction can be determined in real-time based on the specific face being viewed and the individual preferences of each patient.
@@ 3.3 Face Perception with Future Implants
The 100m pixels in PRIMA impose a hard limit on the spatial frequencies that can be perceived by patients. To improve the resolution limit, a new chip with smaller pixels has been tested in rats and has shown to provide higher resolution (wang_electronic_2022). This technology is currently being transferred to Science Corporation for the fabrication of its human version. To illustrate the potential benefits of smaller pixels for face perception, we simulated prosthetic vision with 2x2mm implants composed of 50m and 20m pixels, with corresponding cutoff frequencies of 20 and 50 cycles per image.
As shown in Figure 10, smaller pixels enable significantly higher spatial frequencies, allowing the finer details of the human face, such as the eyes, nose, eyebrows and lips, to be sampled by the device and visualized. However, despite the improvement in spatial resolution, the reduced contrast of prosthetic vision still makes the simulated image harder to perceive compared to the original. Applying inversed contrast filters and facial landmarking prior to the simulation algorithm optimizes the advantage of smaller pixels. Implants with much higher resolution, such as 20m pixels, might require only the inverse contrast filters without facial landmarking, but that will be determined based on individual patients’ preference in the future clinical trials.
3.4 Testing Face and Emotion Recognition on Sighted Individuals
To test the extent of face perception with prosthetic vision and the efficacy of the proposed enhancements, we conducted a study on 16 sighted individuals (9 males and 7 females) drawn from the student body of Stanford University. The participants were positioned in front of a computer screen displaying a question with four faces as the possible answers. For each question, the participants were requested to choose the correct image. The questions were: (1) Which person is the odd one out? (2) Which person is a different gender? and (3) Which face looks happy, sad, surprised, disgusted, angry, confused, fearful, or neutral.
The trial consisted of two phases: first, the images were processed to simulate prosthetic vision with PRIMA without any enhancements, whereas in the second phase, the same images were enhanced by an inversed tone curve and facial landmarking before applying the prosthetic vision flow (similar to the illustrations in Figure 9). Each type of question was repeated 24 times throughout each phase with varying faces. As there were 8 emotions and 24 repetitions, each emotion was displayed 3 times during each phase. Consequently, the accuracy of any given emotion within each phase was either 0, 33.3, 66.6, or 100%. The tone curve applied in both cases to simulate PRIMA vision was the Sigmoid transform, as it seemed to be in better agreement with the descriptions of the patients. A gain of 30 and a shift of 0.2 were chosen to simulate profound contrast sensitivity loss (as illustrated in Figure 6C). The participants had 20 seconds to choose their answer on each screen, and the response time was recorded. Additional 5 participants that didn’t take part in the original trial were recorded performing the same protocol but using the natural, unprocessed images, to obtain a baseline of the average response time for each task.
The easiest task was identifying the odd person out. Even in the first phase, relying on the hair and the general outline of the face, participants demonstrated an average accuracy of 96.1%, with no significant increase in the second phase. The average response time decreased from 5.12 to 4.32 seconds (p<0.001), with the baseline response time for this task being 3.33 seconds. Gender was also identified with a relatively high average accuracy of 76.8% in the first phase, with no significant increase in the second phase. The response time decreased from 6.52 to 5.84 seconds (p<0.05), with the baseline for this task being 4.17 seconds.
As expected, the most challenging task was recognizing emotions. During the first phase, the average accuracy was 25.5% - equivalent to a random choice between four images, indicating the lack of internal facial features essential for emotion recognition. During the second phase, the accuracy increased to 48.4% (p<0.001), demonstrating the benefits of the image enhancements for this task. In addition, the average response time dropped from 8.17 to 6.45 seconds (p<0.001), close to the baseline response time for this task, which was 6 seconds. This outcome is even more significant since some of the presented emotions, such as sad, neutral, and disgusted, were very nuanced. For more obvious emotions, the accuracy of “happy” increased from 35.4% to 83.3% (p<0.001), “confused” - from 25% to 56.25% (p<0.01), “fearful” - from zero to 56.2% (p<0.001), “surprised” – from 37.5% to 64.6% (p<0.05), and “angry” from 35.4% to 56.2% (p<0.05).